OpenStudy (anonymous):

s(x) = sinx Limit (sin(x+h) - sinx)/h siny - sinx = 2sin(y-x)/2 (cos(y+x))/2 Let y=x+h sin(h/2)/(h/2)(cos(x + (h/2)) My problem is I cant figure out how he got the left side of the last equation. I see where the cos(x + (h/2)) is coming from.

OpenStudy (anonymous):

can you post a screen shot?

OpenStudy (anonymous):

I guess I could try. This is a proof of sinx/x = 1

OpenStudy (anonymous):

Let me know if the text is too small.

OpenStudy (anonymous):

yeah that is usually done by something called the "squeeze theorem" have never seen this approach before but i bet your question is algebra based

OpenStudy (anonymous):

It's using trig identities and to be honest I haven't studied much trig

OpenStudy (anonymous):

yeah it is an algebra question you are asking

OpenStudy (anonymous):

the input for cosine in one line is \(\frac{y+x}{2}\) then they wrote put \(y=x+h\) that makes \[\frac{y+x}{2}=\frac{x+h+x}{2}=\frac{2x+h}{2}=x+\frac{h}{2}\] algebra for all of that

OpenStudy (anonymous):

Yeah, I figured out what he was doing with Cos, I just cant figure out how he got sin(h/2)/(h/2)